solve each system of equations by the substitution method
2x-y=-7
4x-3y=-11

Answer by mananth(16946)


2 x + -1 y = -7 .............1
-1 y = -7 + -2 x
/ -1 =
y = 7 + 2 x
4 x + -3 y = 7 .....................2
Substitute y in (2)
4 x + -3 ( 7 + 2 x ) = 7
4 x + -21 + -6 x = 7
4 x -6 x = 28
-2 x = 28
/ -2
x= -14
Plug the value of x in (1)
2 x + -1 y = -7
2 * -14 + -1 y = -7
-28 + -1 y = -7
-1 y = 21
/ -1
y= -21
m.ananth@hotmail.ca